OpenStudy (anonymous):

Write an equivalent expression using radical notation 256^3/4 is this correct????? ^4√256

OpenStudy (jamesj):

So you want \[256^{3/4}\] yes?

OpenStudy (anonymous):

yes please

OpenStudy (jamesj):

Well\[x^{2/3} = \sqrt[3]{x^2}\]

OpenStudy (jamesj):

and in general \[x^{m/n} = (x^m)^{1/n} = \sqrt[n]{x^m}\]

OpenStudy (anonymous):

so am I just missing the 3 ?

OpenStudy (jamesj):

So \[256^{3/4} = (256^3)^{1/4} = ...what?\]

OpenStudy (anonymous):

^4√256^3 ????

OpenStudy (jamesj):

yes